• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 14:22:24 on November 17, 2008

    Entry number 249399

    keyword=Experiment status and plan, and meeting reminder

    Dear collaborators,

    The ARC energy measurement was completed

    http://www.jlab.org/~adaq/halog/html/0811_archive/081117103234.html

    and the value found by Arun is : E = 5889.4 +/- 0.5 (Stat.) +/- 1.0 (Sys.) MeV

    MCC is taking this value into account for finding a Wien angle which reduces less the polarization of Hall A.

    we asked for an access to check the cryo of the spectrometer. Some adjustement of the flow were made by Heidi.

    A few various activities were done opportunistically : - BB missing ADC channel was fixes by Kalyan - Vince checked the retiming and the missing S2m channel - Jack checked the ventilation for the raster. - I adjusted a bit the phase for the RICH

    Controlled access is done. We will check the beam and take a bit more of data.

    Eugene will do a Moeller measurement after to help with the Wien angle setting.

    We will do the compton tune through the chicane right after.

    We will go back to production after on He3 at 12 uA.

    I remind we have the 4PM meeting this afternoon at second house of counting house, to discuss about status and plans for the experiment.